OpenStudy (ishipdestiel):

Help in 4th grade math- i dont know what this is( its for my btother) Decompose 8/12 in two ways ( but the denominator for both has to be 2)

Directrix (directrix):

>( but the denominator for both has to be 2) Did you mean to write 12 rather than 2 in the above statement?

OpenStudy (ishipdestiel):

Nope that is why I was confused when i was helping my little brother

563blackghost (563blackghost):

To simplify the fraction we can divide by 2/2 in which what we get we can divide once more by 2/2 which should get a final answer with 2/3.... \(\Huge{\frac{8}{12} \div \frac{2}{2} \rightarrow \frac{4}{6} \div \frac{2}{2} \rightarrow \color{red}{\frac{2}{3}}}\)

OpenStudy (ishipdestiel):

The answer has to be like something/ 2